OpenStudy (anonymous):

Sarah starts at one corner of a square field and walks 60 feet alone one side to another corner of the field. She turns 90 degrees and walks 40 feet, and then walks straight back to where she started. What is the area of the part of the field she walked?

OpenStudy (anonymous):

walked around lol

OpenStudy (anonymous):

@dannyrod2000 @_Michaeld22

OpenStudy (dannyrod2000):

? I HAVE NO IDEA

OpenStudy (anonymous):

I know this question scared me to lol

OpenStudy (dannyrod2000):

is Moscow the capital of Russia?

OpenStudy (anonymous):

yes it is lol random I mean Im pretty sure it is

OpenStudy (anonymous):

google to make sure

OpenStudy (_michaeld22):

 I've gotta say, it's finding the area of a triangle. A = 1/2bh B = 40 H = 60 60*40 = 2400. \[2400\div2=1200ft ^{2}\]